Problem #4

The inspection of cans of juice involves sampling some cans and weighing them. The overall average for the samples is 44 ounces. Each sample contains 39 cans. The standard deviation is estimated to be 4 ounces. With a 95% confidence interval, develop a control chart with upper and lower control limits. Briefly discuss your findings.
Answer #1

Given, mean = μ = 44

Sample Size = n = 39

Sample Standard Deviation = Population standard deviation / √ n = σ = 4/√ 39

From z table, for 95% confidence interval, z = 1.65

Hence, UCL = μ + zσ = 44 + 1.65*4/√ 39 = 45.06

LCL = μ - zσ = 44 - 1.65*4/√ 39 = 42.94